Session: May 1998 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Maria J. Rioja at Tue Jun 2 08:16:45 METDST 1998) Success
No problems.

Torun (K. Borkowski at Tue Jun 2 10:22:16 METDST 1998) Success
No known problems.

Westerbork (Tony Foley at Tue Jun 2 10:57:33 METDST 1998) Minor failures
Tape ran off the end of reel at 03:58, just like in ef005a, after a fast forward from 7000ft. Hence only 01:00 to 03:58 recorded. I will have to check the low tape sensor in forward.

Onsala (A. Polatidis at Tue Jun 2 14:00:24 METDST 1998) Minor failures
We lost the first scan because the tape lost vacuum. We loaded a new tape (ATNF0022,0707) and started recording from pass 2 (reverse). Although it had been prepassed recently the faulty tape, when inspected after the experiment, appeared loosely packed on the metal reel. There were no problems after the second pass.

Effelsberg (dgraham at Tue Jun 2 18:49:06 METDST 1998) Minor failures
light rain, otherwise OK

Jodrell (Alastair Gunn at Wed Jun 3 10:59:42 METDST 1998) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Mon Jun 8 13:31:22 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name ef005b.antab. Read the comments in the file to get further information.

Noto (C.Trigilio at Mon Jun 8 17:35:23 METDST 1998) Success